# DASPack: Controlled Data Compression for Distributed Acoustic SensingDASPack: Controlled Data Compression for Distributed Acoustic Sensing

### v1

@@ 3 Universitat Politècnica de Catalunya (UPC)
We present DASPack, a high-performance, open-source compression tool specifically designed for distributed acoustic sensing (DAS) data. As DAS becomes a key technology for real-time, high-density, and long-range monitoring in fields such as geophysics, infrastructure surveillance, and environmental sensing, the volume of collected data is rapidly increasing. Large-scale DAS deployments already generate hundreds of terabytes and are expected to increase in the coming years, making long-term storage a major challenge. Despite this urgent need, few compression methods have proven to be both practical and scalable in real-world scenarios. DASPack is a fully operational solution that consistently outperforms existing techniques for DAS data. It enables both controlled lossy and lossless compression by allowing users to choose the maximum absolute difference per datum between the original and compressed data. The compression pipeline combines wavelet transforms, linear predictive coding, and entropy coding to optimise efficiency. Our method achieves up to 3×3$\times$3 × file size reductions for strain and strain rate data in lossless mode across diverse datasets. In lossy mode, compression improves to 6×6$\times$6 × with near-perfect signal fidelity, and up to 10×10$\times$10 × is reached with acceptable signal degradation. It delivers fast throughput (100-200 MB s-1 using a single-thread and up to 750 MB s-1 using 8-threads), enabling real-time deployment even under high data rates. We validated its performance on 15 datasets from a variety of acquisition environments, demonstrating its speed, robustness, and broad applicability. DASPack provides a practical foundation for long-term, sustainable DAS data management in large-scale monitoring networks.
@@ 1 Introduction
Recent studies have explored deep-learning-based compression for DAS. Ni et al. (2024) used a lightweight model to achieve reasonable wavefield reconstruction from 5×5$\times$5 × subsampling. Chen et al. (2025) achieved up to 8×8$\times$8 × lossy compression using a vision transformer on 10-second DAS tiles, which can be further enhanced through downsampling techniques to achieve greater reductions in data size. However, learned compression methods currently face significant limitations. These include a lack of generalisability across datasets, inability to guarantee explicit error bounds, and the tendency to introduce data-driven reconstruction artifacts that can inadvertently regenerate patterns present in the training data. Moreover, such models are computationally intensive, often requiring resources that prevent real-time deployment, especially on CPU-based systems. Only lightweight variants with moderate compression power can operate at reasonable speed. As a result, these approaches remain impractical for many real-world scenarios in their current form.
@@ 3.2 Controlled degradation
Unlike filtering or decimation, quantisation preserves the original signal bandwidth, maintaining both temporal and spatial resolution. For smaller $\Delta$, the rounding error tends to be uniform and uncorrelated across time and channels, resembling additive white noise.
@@ 3.3 Compression Pipeline
    Then, a discrete wavelet transform is optionally applied along both the time and distance dimensions (Fig. 2c). The wavelet transform decomposes data in several frequency and wavenumber bands, without increasing the total number of samples and allowing afterward a perfect reconstruction of the original signal.
@@ 4.1 Compression efficiency and speed
Figs. 6 and 7 present a comparative analysis of the trade-off between compression speed (throughput in MB s-1 ) and efficiency (compression factor) for five compression methods (gzip, zstd, JPEG2000, H5TurboPFor, and DASPack) at different accuracies. Experiments were run in a conventional desktop computer (Apple M1 Pro CPU) in single-thread mode. Fig. 6 focuses on strain rate datasets 1-8, all recorded as 16-bit integers, while Fig. 7 focuses on strain datasets (9-15), recorded as 32-bit integers or floating-points. Because H5TurboPFor only works on 16-bit data, 32-bit integers are interpreted as two 16-bit values for this particular method. We fixed DASPack’s parameters to square tiles of 2000 samples per side and two LPC coefficients, and implemented the steps using 32-bit integer operations. Strain data generally spans a higher dynamic range than strain rate due to the presence of high-amplitude, long period signals, hence requiring more bits to store raw data. Yet, for this same reason, strain data exhibit higher temporal correlation, hence achieving higher compression ratios.
Figures 6a and 7a reveal the performance of different methods for each individual dataset at a given fixed-accuracy (lossless for strain rate, $\Delta=10^{-1.5}~{}\text{n}\varepsilon$ for strain). Overall, gzip and zstd achieve moderate compression factors but exhibit relatively low throughput, generally below 100 MB/s. H5TurboPFor and Zfp prioritise speed, often exceeding 200 MB/s, and H5TurboPFor maintains high compression factors for strain rate data. Zfp performs poorly in strain rate because it is a method optimised for floating-points; but the performance is only moderately better for strain data. JPEG2000 achieves higher compression factor, but is the slowest method, achieving 10-20 MB s-1. DASPack, by contrast, offers both high throughput (up to 250 MB s-1 for strain, 125 MB s-1 for strain rate) and the strongest compression performance. Performance varies by dataset, highlighting the importance of balancing speed and efficiency depending on data characteristics and application needs. DASPack consistently occupies a favourable region of the plot, offering high compression ratios at speeds that match or exceed leading alternatives. This makes it a strong choice for applications requiring both data reduction and fast, scalable performance.
@@
JPEG2000 and DASPack provide consistently superior compression performance at higher tolerances, reaching compression factors beyond 8×8$\times$8 ×. Even more relaxed accuracies (such as $1~{}\text{n}\varepsilon$ or a $10^{2}$ quantisation step) can result in up to 10×10$\times$10 × compression for some datasets. In certain applications or datasets, this may offer a viable trade-off without compromising critical signal content.
@@ 4.2 Application examples
Fig. 9 presents filtered strain data from a magnitude mbLg 3.1 earthquake recorded in the Alboran Sea (dataset 10). This example illustrates how DASPack performs when the signal amplitude is very small (only about ±$\pm 2.5~{}\text{n}\varepsilon$ across the entire event). In this case, quantisation steps of $\Delta=0.1~{}\text{n}\varepsilon$ and $\Delta=1~{}\text{n}\varepsilon$ result in compression factors of $\times6$ and 11.9×11.9$\times$11.9 ×, respectively. At the finer step, the quantised waveform closely matches the original, with negligible error. At $1~{}\text{n}\varepsilon$, the error map shows residuals well below $0.5~{}\text{n}\varepsilon$, indicating that most of the signal energy survives even at this aggressive compression level.
@@
Similar patterns are observed in Fig. 10, which shows filtered strain data from a quarry blast recorded during the Vinaroz experiment (dataset 13), with amplitudes reaching about ±$\pm 10~{}\text{n}\varepsilon$. Quantisation steps of $\Delta=0.1~{}\text{n}\varepsilon$ and $\Delta=1~{}\text{n}\varepsilon$ result in compression factors of 5.2×5.2$\times$5.2 × and 9.0×9.0$\times$9.0 ×, respectively. In this example, the error is not distributed uniformly: offshore channels (beyond 10 km) retain nearly all signal information even at the higher quantisation level, while the onshore section (within the first 10 km) records lower amplitudes and threfore exhibits higher residuals in the error maps.
@@
Finally, Fig. 11 extends the examples to a high-amplitude (±$\pm 100~{}\text{n}\varepsilon$), 4.9 mbLg earthquake recorded during the 2021 La Palma eruption (dataset 15). This case confirms that DASPack maintains high fidelity at moderate quantisation. For example, $\Delta=4~{}\text{n}\varepsilon$ results in a compression factor of 9.1×9.1$\times$9.1 ×.
@@
Figure 8: A regional earthquake of magnitude ML 3.4 occurred 48 km northwest of the interrogator on 23 September 2021 at 07:06:54 UTC during the Athens experiment (dataset 2). The figure compares unfiltered raw signals with quantised versions using two degradation levels ($\Delta=2~{}\text{n}\varepsilon~{}\text{s}^{-1}$ and $\Delta=8~{}\text{n}\varepsilon~{}\text{s}^{-1}$), corresponding to compression factors of $c_{f}=5.5\times$ and $c_{f}=9.2\times$. (a) Waveform from channel 6000 (at 6000 m). (b) Zoomed view of the waveform. (c) 2-D strain map of the original signal (left) and error maps between the original and degraded signals (middle and right).
@@
Figure 9: An onshore earthquake of magnitude mbLg 3.1 occurred 97 km northeast of the interrogator on 5 October 2023 at 21:40:04 UTC and was recorded on a submarine fibre-optic cable, in the Alboran Sea (dataset 10). The figure compares filtered raw signals with quantised versions using two degradation levels ($\Delta=0.1~{}\text{n}\varepsilon$ and $\Delta=1~{}\text{n}\varepsilon$), corresponding to compression factors of $c_{f}=6.6\times$ and $c_{f}=11.9\times$. (a) Waveform from channel 1700 (at 17,000 m). (b) Zoomed view of the waveform. (c) 2-D strain map of the original signal (left) and error maps between the original and degraded signals (middle and right).
@@
Figure 10: A quarry blast occurred 26 km southeast of the interrogator on 9 June 2023 at 10:26:30 UTC and was recorded during the Vinaroz experiment (dataset 13). The figure compares filtered raw signals with quantised versions using two degradation levels ($\Delta=0.1~{}\text{n}\varepsilon$ and $\Delta=1~{}\text{n}\varepsilon$), corresponding to compression factors of $c_{f}=5.2\times$ and $c_{f}=9.0\times$. (a) Waveform from channel 900 (at 9,000 m). (b) Zoomed view of the waveform. (c) 2-D strain map of the original signal (left) and error maps between the original and degraded signals (middle and right).
@@
Figure 11: An earthquake of magnitude mbLg 4.9 occurred 148 km west from the interrogator on 26 October 2021 at 16:25:37 UTC during the La Palma volcanic eruption and was recorded at the Teide Astronomical Observatory (dataset 14). The figure compares filtered raw signals with quantised versions using two degradation levels ($\Delta=0.1~{}\text{n}\varepsilon$ and $\Delta=4~{}\text{n}\varepsilon$), corresponding to compression factors of $c_{f}=4.1\times$ and $c_{f}=9.1\times$. (a) Waveform from channel 1000 (at 10,000 m). (b) Zoomed view of the waveform. (c) 2-D strain map of the original signal (left) and error maps between the original and degraded signals (middle and right).
@@ 5.1 Compression efficiency
DASPack contributes to DAS data management by providing sustained, real-time throughput while simultaneously delivering the highest compression factors across a broad range of acquisition settings. Its performance advantage is most pronounced for strain data, where the combination of wavelet decomposition and linear predictive coding (LPC) removes the slowly varying, large-amplitude components that dominate the signal variance. Because strain is typically stored in 32-bit words to accommodate this broad dynamic range, both the entropy reduction and the arithmetic-coding efficiency are amplified, yielding compression factors that exceed 8×8$\times$8 ×. strain rate, in contrast, is acquired with narrower dynamic range and usually quantised at 16 bits; here the algorithm still outperforms specialised competitors with smaller margins, confirming that the same pipeline remains effective even when the predictable, low-frequency portion of the spectrum is reduced. This difference in performance arises not from the compression method itself but from the statistical properties of the input data: as a temporal derivative of strain, strain rate naturally exhibits lower amplitude variability and reduced temporal coherence, which limits the predictive power of both wavelet and LPC stages and reduces overall compressibility.
@@ 5.2 Compression speed and real-time capability
In practice, DASPack uses 32-bit integer operations and achieves a sustained compression speed of approximately 250 MB s-1 for strain data when running in single-threaded mode. For strain rate data, typically stored as 16-bit integers, performance is about half, around 125 MB s-1. Thanks to independent per-tile processing, DASPack supports parallelism. hanks to its tile-based processing design, DASPack supports parallel execution. When configured with 8 threads on an Apple M1 Pro CPU, the average compression speed across our datasets increases to roughly 750 MB s-1. This throughput exceeds the write speed of standard hard disk drives (HDDs), which typically peak around 200 MB/s, making storage, rather than computing, the limiting factor in many setups. To fully leverage DASPack’s compression speed, faster storage such as solid-state drives (SSDs), which can sustain several gigabytes per second, is recommended.

### v2

@@ 3 Universitat Politècnica de Catalunya (UPC)
We present DASPack, a high-performance, open-source compression tool specifically designed for distributed acoustic sensing (DAS) data. As DAS becomes a key technology for real-time, high-density, and long-range monitoring in fields such as geophysics, infrastructure surveillance, and environmental sensing, the volume of collected data is rapidly increasing. Large-scale DAS deployments already generate hundreds of terabytes and are expected to increase in the coming years, making long-term storage a major challenge. Despite this urgent need, few compression methods have proven to be both practical and scalable in real-world scenarios. DASPack is a fully operational solution that consistently outperforms existing techniques for DAS data. It enables both controlled lossy and lossless compression by allowing users to choose the maximum absolute difference per datum between the original and compressed data. The compression pipeline combines wavelet transforms, linear predictive coding, and entropy coding to optimise efficiency. Our method achieves up to $3\times$ file size reductions for strain and strain rate data in lossless mode across diverse datasets. In lossy mode, compression improves to $6\times$ with near-perfect signal fidelity, and up to $10\times$ is reached with acceptable signal degradation. It delivers fast throughput (100–200 MB s-1 using a single-thread and up to 750 MB s-1 using 8-threads), enabling real-time deployment even under high data rates. We validated its performance on 15 datasets from a variety of acquisition environments, demonstrating its speed, robustness, and broad applicability. DASPack provides a practical foundation for long-term, sustainable DAS data management in large-scale monitoring networks.
@@ 1 Introduction
Recent studies have explored deep-learning-based compression for DAS. Ni et al. (2024) used a lightweight model to achieve reasonable wavefield reconstruction from $5\times$ subsampling. Chen et al. (2025) achieved up to $8\times$ lossy compression using a vision transformer on 10-second DAS tiles, which can be further enhanced through downsampling techniques to achieve greater reductions in data size. However, learned compression methods currently face significant limitations. These include a lack of generalisability across datasets, inability to guarantee explicit error bounds, and the tendency to introduce data-driven reconstruction artifacts that can inadvertently regenerate patterns present in the training data. Moreover, such models are computationally intensive, often requiring resources that prevent real-time deployment, especially on CPU-based systems. Only lightweight variants with moderate compression power can operate at reasonable speed. As a result, these approaches remain impractical for many real-world scenarios in their current form.
@@ 3.2 Controlled degradation
Unlike filtering or decimation, quantisation preserves the original signal bandwidth, maintaining both temporal and spatial resolution. For smaller $\Delta$, the rounding error tends to be uniform and uncorrelated across time and channels, resembling additive white noise (Widrow & Kollár, 2008).
@@ 3.3 Compression Pipeline
    Then, a discrete wavelet transform is optionally applied along both the time and distance dimensions (Fig. 2c). The wavelet transform decomposes data in several frequency and wavenumber bands along the time and distance dimensions, respectively, without increasing the total number of samples and allowing afterward a perfect reconstruction of the original signal (Jacques et al., 2011).
@@ 4.1 Compression efficiency and speed
Figs. 6 and 7 present a comparative analysis of the trade-off between compression speed (throughput in MB s-1 ) and efficiency (compression factor) for five compression methods (gzip, zstd, JPEG2000, H5TurboPFor, Zfp and DASPack) at different accuracies. Experiments were run in a conventional desktop computer (Apple M1 Pro CPU) in single-thread mode. Fig. 6 focuses on strain rate datasets 1-8, all recorded as 16-bit integers, while Fig. 7 focuses on strain datasets (9-15), recorded as 32-bit integers or floating-points. Because H5TurboPFor only works on 16-bit data, 32-bit integers are interpreted as two 16-bit values for this particular method. We fixed DASPack’s parameters to square tiles of 2000 samples per side and two LPC coefficients, and implemented the steps using 32-bit integer operations. Strain data generally spans a higher dynamic range than strain rate due to the presence of high-amplitude, long period signals, hence requiring more bits to store raw data. Yet, for this same reason, strain data exhibit higher temporal correlation, hence achieving higher compression ratios.
Figures 6a and 7a reveal the performance of different methods for each individual dataset at a given fixed-accuracy (lossless for strain rate, $\Delta=10^{-1.5}~\text{n}\varepsilon$ for strain). Overall, gzip and zstd achieve moderate compression factors but exhibit relatively low throughput, generally below 100 MB/s. H5TurboPFor and Zfp prioritise speed, often exceeding 200 MB/s, and H5TurboPFor maintains high compression factors for strain rate data. Zfp performs poorly in strain rate because it is a method optimised for floating-points; but the performance is only moderately better for strain data. JPEG2000 achieves higher compression factor, but is the slowest method, achieving 10-20 MB s-1. DASPack, by contrast, offers both high throughput (up to 250 MB s-1 for strain, 125 MB s-1 for strain rate) and the strongest compression performance. It is important to note that compression factors are dataset-dependent and should not be taken as absolute numbers. While different datasets may yield higher or lower compression factors, the consistent trend is that DASPack outperforms alternative methods across all cases considered. DASPack consistently occupies a favourable region of the plot, offering high compression ratios at speeds that match or exceed leading alternatives. This makes it a strong choice for applications requiring both data reduction and fast, scalable performance.
@@
JPEG2000 and DASPack provide consistently superior compression performance at higher tolerances, reaching compression factors beyond $8\times$. Even more relaxed accuracies (such as $1~\text{n}\varepsilon$ or a $10^{2}$ quantisation step) can result in up to $10\times$ compression for some datasets. In certain applications or datasets, this may offer a viable trade-off without compromising critical signal content.
@@ 4.2 Application examples
Fig. 9 presents filtered strain data from a magnitude mbLg 3.1 earthquake recorded in the Alboran Sea (dataset 10). This example illustrates how DASPack performs when the signal amplitude is very small (only about $\pm 2.5~\text{n}\varepsilon$ across the entire event). In this case, quantisation steps of $\Delta=0.1~\text{n}\varepsilon$ and $\Delta=1~\text{n}\varepsilon$ result in compression factors of $6.6\times$ and $14.4\times$, respectively. At the finer step, the quantised waveform closely matches the original, with negligible error. At $1~\text{n}\varepsilon$, the error map shows residuals below $0.5~\text{n}\varepsilon$ that do not exhibit any visual patterns, indicating that most of the signal energy survives even at this aggressive compression level.
@@
Similar patterns are observed in Fig. 10, which shows filtered strain data from a quarry blast recorded during the Vinaroz experiment (dataset 13), with amplitudes reaching about $\pm 10~\text{n}\varepsilon$. Quantisation steps of $\Delta=0.1~\text{n}\varepsilon$ and $\Delta=1~\text{n}\varepsilon$ result in compression factors of $5.7\times$ and $11.6\times$, respectively. In this example, the error is not distributed uniformly: offshore channels (beyond 10 km) retain nearly all signal information even at the higher quantisation level, while the onshore section (within the first 10 km) records lower amplitudes and threfore exhibits higher residuals in the error maps.
@@
Finally, Fig. 11 extends the examples to a high-amplitude ($\pm 100~\text{n}\varepsilon$), 4.9 mbLg earthquake recorded during the 2021 La Palma eruption (dataset 15). This case confirms that DASPack maintains high fidelity at moderate quantisation. For example, $\Delta=4~\text{n}\varepsilon$ results in a compression factor of $10.9\times$. In this higher degradation case, the error appears to be mostly uncorrelated, with higher values along the high-amplitude front.
@@
Figure 8: A regional earthquake of magnitude ML 3.4 occurred 48 km northwest of the interrogator on 23 September 2021 at 07:06:54 UTC during the Athens experiment (dataset 2). The figure compares unfiltered raw signals with quantised versions using two degradation levels ($\Delta=2~\text{n}\varepsilon~\text{s}^{-1}$ and $\Delta=8~\text{n}\varepsilon~\text{s}^{-1}$), corresponding to compression factors of $c_{f}=6.1\times$ and $c_{f}=9.7\times$. (a) Waveform from channel 6000 (at 6000 m). (b) Zoomed view of the waveform. (c) Discrete Fourier Transform of the zoomed window. (d) 2-D strain map of the original signal (left) and error maps between the original and degraded signals (middle and right).
@@
Figure 9: An onshore earthquake of magnitude mbLg 3.1 occurred 97 km northeast of the interrogator on 5 October 2023 at 21:40:04 UTC and was recorded on a submarine fibre-optic cable, in the Alboran Sea (dataset 10). The figure compares filtered raw signals with quantised versions using two degradation levels ($\Delta=0.1~\text{n}\varepsilon$ and $\Delta=1~\text{n}\varepsilon$), corresponding to compression factors of $c_{f}=6.6\times$ and $c_{f}=14.4\times$. (a) Waveform from channel 1700 (at 17,000 m). (b) Zoomed view of the waveform. (c) Discrete Fourier Transform of the zoomed window. (d) 2-D strain map of the original signal (left) and error maps between the original and degraded signals (middle and right).
@@
Figure 10: A quarry blast occurred 26 km southeast of the interrogator on 9 June 2023 at 10:26:30 UTC and was recorded during the Vinaroz experiment (dataset 13). The figure compares filtered raw signals with quantised versions using two degradation levels ($\Delta=0.1~\text{n}\varepsilon$ and $\Delta=1~\text{n}\varepsilon$), corresponding to compression factors of $c_{f}=5.7\times$ and $c_{f}=11.6\times$. (a) Waveform from channel 900 (at 9,000 m). (b) Zoomed view of the waveform. (c) Discrete Fourier Transform of the zoomed window. (d) 2-D strain map of the original signal (left) and error maps between the original and degraded signals (middle and right).
@@
Figure 11: An earthquake of magnitude mbLg 4.9 occurred 148 km west from the interrogator on 26 October 2021 at 16:25:37 UTC during the La Palma volcanic eruption and was recorded at the Teide Astronomical Observatory (dataset 14). The figure compares filtered raw signals with quantised versions using two degradation levels ($\Delta=0.1~\text{n}\varepsilon$ and $\Delta=4~\text{n}\varepsilon$), corresponding to compression factors of $c_{f}=6.1\times$ and $c_{f}=10.9\times$. (a) Waveform from channel 1000 (at 10,000 m). (b) Zoomed view of the waveform. (c) Discrete Fourier Transform of the zoomed window. (d) 2-D strain map of the original signal (left) and error maps between the original and degraded signals (middle and right).
@@ 5.1 Compression efficiency
DASPack contributes to DAS data management by providing sustained, real-time throughput while simultaneously delivering the highest compression factors across a broad range of acquisition settings. Its performance advantage is most pronounced for strain data, where the combination of wavelet decomposition and linear predictive coding (LPC) removes the slowly varying, large-amplitude components that dominate the signal variance. Because strain is typically stored in 32-bit words to accommodate this broad dynamic range, both the entropy reduction and the arithmetic-coding efficiency are amplified, yielding compression factors that exceed $8\times$. strain rate, in contrast, is acquired with narrower dynamic range and usually quantised at 16 bits; here the algorithm still outperforms specialised competitors with smaller margins, confirming that the same pipeline remains effective even when the predictable, low-frequency portion of the spectrum is reduced. This difference in performance arises not from the compression method itself but from the statistical properties of the input data: as a temporal derivative of strain, strain rate naturally exhibits lower amplitude variability and reduced temporal coherence, which limits the predictive power of both wavelet and LPC stages and reduces overall compressibility.
@@ 5.2 Compression speed and real-time capability
In practice, DASPack uses 32-bit integer operations and achieves a sustained compression speed of approximately 250 MB s-1 for strain data when running in single-threaded mode. For strain rate data, typically stored as 16-bit integers, performance is about half, around 125 MB s-1. Thanks to independent per-tile processing, DASPack supports parallelism. When configured with 8 threads on an Apple M1 Pro CPU, the average compression speed across our datasets increases to roughly 750 MB s-1. This throughput exceeds the write speed of standard hard disk drives (HDDs), which typically peak around 200 MB s-1, making storage, rather than computing, the limiting factor in many setups. To fully leverage DASPack’s compression speed, faster storage such as solid-state drives (SSDs), which can sustain several gigabytes per second, is recommended. Clusters with a large number of nodes and drive bays offering high read and write speeds further exploit parallelism, enabling highly efficient compression and decompression.
5.3 Understanding the effect of quantisation
The controlled degradation step in DASPack is based on fixed-accuracy quantisation, which has well-understood statistical properties. For sufficiently small quantisation steps, the rounding error behaves like additive uniform noise with zero mean and known variance (Widrow & Kollár, 2008). This behaviour explains why the resulting error is often modeled as white noise, and was experimentally observed in Figures 811b,d. However, unlike truly independent noise, quantisation can lead to repeated values in low-amplitude segments. For example, entire portions of a trace being rounded to zero. At very coarse quantisation levels, this may introduce visible artifacts, even if the overall frequency content remains largely preserved.
Figures 811c illustrate that the frequency spectrum is barely modified by quantisation, and only the very low-amplitude parts of the signal are suppressed. Indeed, interpreted as random noise, the quantisation error converges to Gaussian behaviour in the frequency domain (Oppenheim & Schafer, 2009).
Another perspective on quantisation is to view it as a uniform degradation of the signal-to-noise ratio (SNR). If the error behaves like additive uniform noise, its power increases proportionally to the square of the quantisation step (Widrow & Kollár, 2008). This is consistent with the results in Figure 13, where the SNR was computed as the ratio of signal and noise window powers. To prevent spurious vanishing errors caused by large quantisation steps rounding entire traces to zero, dithering was added during the SNR calculation, which helps recover idealised behaviour of quantisation noise (Widrow & Kollár, 2008). Importantly, dithering was introduced solely during SNR evaluation to obtain meaningful estimates at coarse quantisation levels; it is not part of the DASPack method itself. The measured SNR follows closely the expected noise law, providing an additional validation of the theoretical model. This analysis demonstrates that quantisation in DASPack is not only interpretable as a bounded reconstruction error, but can also be understood within the familiar framework of additive noise, with clear implications for the effective resolution and dynamic range of compressed DAS signals.
While the statistical interpretation of quantisation is useful, it does not by itself capture the full implications for downstream seismic analysis. DAS is applied to a wide variety of problems, each of which may respond differently to small compression errors. For instance, typical processing tasks include (i) detecting an event through STA/LTA or template matching, (ii) picking traveltimes using manual, automatic, or array-based approaches, (iii) estimating dispersion curves from ambient noise correlations, or (iv) measuring amplitude decay to infer attenuation. Each of these workflows involves distinct sensitivities to low-amplitude signals, phase accuracy, or spectral content. A superficial treatment of all these aspects would risk being incomplete or even misleading. For this reason, we deliberately refrain from providing application-specific guidelines in this paper. Instead, we emphasise that a rigorous evaluation of the effect of controlled degradation on the accuracy of diverse seismological analyses deserves a dedicated study.
Figure 13: Signal-to-noise ratio (SNR) as a function of quantisation step $\Delta$ for two DAS datasets. (a) La Palma eruption (26 October 2021, channel #1000), also shown in Fig. 11 and (b) Athens array (23 September 2021, channel #6000), also shown in Fig. 8. Grey markers denote measured SNR values, while dashed green lines show fitted theoretical predictions of the form $10\log_{10}\frac{P_{s}}{P_{n}+k\Delta^{2}}$. The fitted constants are $k=0.013$ ($R^{2}=0.95$) for La Palma and $k=0.034$ ($R^{2}=0.87$) for Athens. For comparison, ideal i.i.d. uniform quantisation noise would correspond to $k=1/12\approx 0.083$. Dithering was applied during the SNR calculation to mitigate artificial vanishing errors at large $\Delta$.